How much of a 20% acid solution should be added to 10 L of a 60% acid solution to make a 25% acid solution?
Answer by mananth(16946)

--- percent ---------------- Amount
Solution I 60 ---------------- 10 liters
solution II 20 ---------------- x liters
Mixture 25 ---------------- 10 + x liters

60 * 10 + 20 x = 25 ( 10 + x )
600 + 20 x = 250 + 25 x
20 x -25 x = -600 + 250
-5 x = -350
/ -5
x = 70 liters solution II

m.ananth@hotmail.ca